### v1

@@ 2.4 Prediction for Different Clinical Endpoints Trials
The figure (Fig.1 C) illustrates the distribution of clinical trial outcomes across different model datasets(Table 4). Overall Survival (OS) dominates the outcome categories for all models, consistently representing around 64-70% of the data. Objective Response Rate (ORR) follows as the second most represented category, ranging from 6% to 11%. Progression-Free Survival (PFS) is included in a smaller proportion, around 3-4% across models. Safety indicators (AE and SAE) occupy a small portion, varying between 1.6% and 4.9%. Lastly, the ”Other” category (for example: Relapse-Free Survival, Pharmacokinetic Parameters, Patient-Reported Outcomes) accounts for around 14-19% of the outcomes across the datasets. Most clinical trials prioritize Overall Survival (OS) as the primary endpoint, and model performance on OS closely mirrors their balanced accuracy and MCC values.
@@
For other models, performance across different clinical trial endpoint categories showed similar trends. For instance, among all models that exhibited some recognition capability, the ability to identify clinical trials with Overall Survival (OS) as the endpoint was the strongest. For GPT-4o and GPT-4, balanced accuracy in this category was significantly above the average (0.614 and 0.562). This phenomenon was also observed in HINT and Llama3.
Safety indicators (AE and SAE) also demonstrated some recognition capabilities in these models. Although performance in this category was not as strong as in the OS category, metrics for GPT-4, GPT-4o, HINT, and Llama3 were all above random classification. In particular, GPT-4o’s specificity in this category reached 0.5, marking the highest level achieved by large language models in this experiment.
@@
For other clinical trial classifications, such as Objective Response Rate (ORR), Progression-Free Survival (PFS), and other uncategorized endpoints, most of these models did not exhibit classification capabilities. Metrics for these models remained at random classification levels. Notably, for GPT-4o, which typically performed well, the MCC for ORR was -0.167, even below 0. Similarly, the usually stable HINT model showed MCC values of -0.022 and -0.12 in these two categories.
@@
Nonetheless, the specificity of the HINT model remained much higher than that of the large language models. This suggests that HINT still maintained a relatively strong ability to recognize negative cases in these categories.
| Model | Endpoint | Trials Number | Balanced_acc | MCC | Recall | Specificity |
| Model: gpt4mini | ||||||
| Overall Survival (OS) | 28 | 0.500 | 0.0 | 1.000 | 0.0 | |
| Objective Response Rate (ORR) | 10 | 0.500 | 0.0 | 1.000 | 0.0 | |
| Progression-Free Survival (PFS) | 5 | 0.500 | 0.0 | 1.000 | 0.0 | |
| Safety Indicators (AE and SAE) | 12 | 0.500 | 0.0 | 1.000 | 0.0 | |
| Others | 13 | 0.000 | 0.0 | 1.000 | 0.0 | |
| Model: gpt4o | ||||||
| Overall Survival (OS) | 28 | 0.614 | 0.278 | 1.000 | 0.333 | |
| Objective Response Rate (ORR) | 10 | 0.438 | -0.167 | 0.968 | 0.438 | |
| Progression-Free Survival (PFS) | 5 | 0.500 | 0.000 | 0.946 | 0.0 | |
| Safety Indicators (AE and SAE) | 12 | 0.583 | 0.169 | 0.962 | 0.500 | |
| Others | 13 | 0.517 | 0.000 | 0.962 | 0.158 | |
| Model: gpt4 | ||||||
| Overall Survival (OS) | 20 | 0.562 | 0.281 | 1.000 | 0.125 | |
| Objective Response Rate (ORR) | 9 | 0.500 | 0.000 | 1.000 | 0.000 | |
| Progression-Free Survival (PFS) | 5 | 0.500 | 0.000 | 1.000 | 0.0 | |
| Safety Indicators (AE and SAE) | 9 | 0.600 | 0.316 | 1.000 | 0.200 | |
| Others | 7 | 1.000 | 0.000 | 1.000 | 0.062 | |
| Model: gpt3.5 | ||||||
| Overall Survival (OS) | 1423 | 0.504 | 0.044 | 1.000 | 0.013 | |
| Objective Response Rate (ORR) | 409 | 0.499 | -0.023 | 0.998 | 0.000 | |
| Progression-Free Survival (PFS) | 163 | 0.506 | 0.040 | 0.998 | 0.029 | |
| Safety Indicators (AE and SAE) | 614 | 0.500 | -0.007 | 0.998 | 0.004 | |
| Others | 463 | 0.504 | 0.000 | 0.998 | 0.012 | |
| Model: HINT | ||||||
| Overall Survival (OS) | 252 | 0.548 | 0.085 | 0.596 | 0.515 | |
| Objective Response Rate (ORR) | 74 | 0.488 | -0.022 | 0.616 | 0.565 | |
| Progression-Free Survival (PFS) | 38 | 0.446 | -0.120 | 0.619 | 0.692 | |
| Safety Indicators (AE and SAE) | 89 | 0.516 | 0.034 | 0.605 | 0.366 | |
| Others | 67 | 0.511 | 0.000 | 0.605 | 0.566 | |
| Model: llama | ||||||
| Overall Survival (OS) | 205 | 0.545 | 0.152 | 0.935 | 0.133 | |
| Objective Response Rate (ORR) | 61 | 0.507 | 0.018 | 0.952 | 0.111 | |
| Progression-Free Survival (PFS) | 26 | 0.476 | -0.098 | 0.951 | 0.000 | |
| Safety Indicators (AE and SAE) | 88 | 0.528 | 0.081 | 0.954 | 0.161 | |
| Others | 45 | 0.507 | 0.000 | 0.954 | 0.062 | |
@@ 2.5 Prediction on different clinical trial period
We observe that for all models, balanced accuracy declines as clinical trial duration increases. Even GPT-4o, which generally performs best among the large language models, sees its balanced accuracy drop from 0.875 for shorter trials to 0.588 for longer trials, a performance barely above random chance. The typically stable HINT model experiences a more drastic decline, with balanced accuracy falling from 0.705 to 0.300, indicating poor prediction capability on extended-duration trials. Its performance on medium-length trials also only marginally exceeds random prediction levels(Table 5).
When examining specificity, the HINT model again shows a declining trend over time, indicating decreased ability to recognize negative cases in longer trials. In contrast, GPT-4, GPT-4o, and other large language models either improve or remain consistent in specificity, demonstrating better or stable performance in recognizing negative samples over longer trials.
This trend suggests that while all models face challenges in predicting outcomes for trials with extended durations, the HINT model struggles particularly in adapting to long-term trials. For these cases, GPT-4 and GPT-4o may offer better positive class identification and more reliable negative class identification in shorter trials, showing relatively higher balanced accuracy and MCC.
| Model | Trial Period | Balanced Accuracy | MCC | Recall | Specificity |
| Model: gpt4mini | |||||
| short-term trials | 1.000 | 0.000 | 1.000 | 0.000 | |
| medium-length trials | 0.500 | 0.000 | 1.000 | 0.000 | |
| long-term trials | 0.500 | 0.000 | 1.000 | 0.000 | |
| Model: gpt4o | |||||
| short-term trials | 0.875 | 0.000 | 0.875 | 0.000 | |
| medium-length trials | 0.625 | 0.316 | 0.917 | 0.333 | |
| long-term trials | 0.562 | 0.265 | 1.000 | 0.125 | |
| Model: gpt4 | |||||
| short-term trials | 1.000 | 0.000 | 1.000 | 0.000 | |
| medium-length trials | 0.500 | 0.000 | 1.000 | 0.000 | |
| long-term trials | 0.583 | 0.320 | 1.000 | 0.167 | |
| Model: gpt3.5 | |||||
| short-term trials | 0.509 | 0.087 | 0.997 | 0.020 | |
| medium-length trials | 0.502 | 0.029 | 0.996 | 0.008 | |
| long-term trials | 0.500 | 0.000 | 1.000 | 0.000 | |
| Model: HINT | |||||
| short-term trials | 0.675 | 0.285 | 0.721 | 0.630 | |
| medium-length trials | 0.519 | 0.034 | 0.521 | 0.517 | |
| long-term trials | 0.333 | -0.342 | 0.211 | 0.455 | |
| Model: llama | |||||
| short-term trials | 0.566 | 0.215 | 0.966 | 0.167 | |
| medium-length trials | 0.483 | -0.065 | 0.942 | 0.024 | |
| long-term trials | 0.548 | 0.145 | 0.929 | 0.167 | |
@@ 2.6 Terminated trials and prediction of termination
Large language models (such as the GPT series and Llama) performed poorly in predicting outcomes for these trials, with accuracy around 0.1 or lower, indicating their difficulty in handling the complex external factors causing trial termination(Table 6). Consequently, we excluded “terminated” trials from further analyses to avoid biasing the experimental results. While large language models could not identify trial termination, the HINT model demonstrated this ability, with stable accuracy across all clinical trial statuses, similar to its performance on the entire dataset, showcasing its effectiveness in distinguishing terminated trials.
| Model | HINT | GPT3.5 | GPT4o | GPT4mini | GPT4 | Llama3 |
@@ 3.3 Clinical Trials’ Label
To label the outcomes of clinical trials, we employed the GPT-4 model to analyze and annotate the trial outcome data[23]. GPT-4 was tasked with classifying each trial as either ”Success” or ”Failure” based on the provided primary and secondary outcome information. Subsequently, human experts reviewed these labels by cross-referencing them with the actual trial outcomes to ensure accuracy. In subsequent analyses, we excluded data with a ” terminated ” status because we found that the LLMs lacked the ability to accurately recognize trial termination. Figure(Fig2 C) illustrates the proportion of different labels across various trial statuses within the entire dataset.
@@ 3.4 Clinical Trials’ Prediction
For the HINT model, we input data in the required format [9], including the SMILES[24] sequence information of the drug and the ICD-10 code of the studied disease(Fig.2 E). Following the preprocessing steps outlined by Fu et al.[bib29], we categorized the data based on the trial phase before feeding it into the HINT model for prediction. The model output is a probability of trial success; results with a success probability greater than 0.5 were labeled as “Success,” while those with a probability less than or equal to 0.5 were labeled as “Failure”.
@@ 3.6 Metrics Selected
Recall (Sensitivity): The proportion of true positives correctly identified among all actual positives. This metric measures the model’s ability to capture relevant instances and avoid false negatives.
@@
| GPT 3.5 | gpt-3.5-turbo-0125 | Sep 2021 |
| GPT 4 | gpt-4-turbo-2024-04-09 | Dec 2023 |
| GPT 4o | chatgpt-4o-latest | Oct 2023 |
| GPT 4mini | gpt-4o-mini-2024-07-18 | Oct 2023 |
| Llama3.1-8B | Llama-3-8B-Instruct | March 2023 |
@@ 4 Conclusion
This study focuses on evaluating the performance of large language models (LLMs) in predicting clinical trial outcomes, with particular attention to Balanced Accuracy, Matthews Correlation Coefficient (MCC), Recall, and Specificity. Among all models, GPT-4o demonstrated the best overall performance, showing high balanced accuracy and MCC, especially in the early trial phases. However, similar to other LLMs, GPT-4o faced limitations in recognizing negative samples, a common challenge across these models.
@@
One significant finding was that all LLMs struggled to predict outcomes for oncology trials, which made up the majority of the dataset. The complexity of cancer-related trials likely contributed to the models’ limited performance in this area. Although LLMs showed strengths in predicting certain outcomes, their limitations in handling complex domains such as oncology constrained their overall predictive capability.
In contrast, the HINT model displayed relatively stable performance across trial phases and showed notable strength in negative sample recognition. Its effectiveness improved with later phases, with Phase 3 showing especially high specificity and strong ability to recognize failed trials. Additionally, HINT demonstrated resilience to external factors, such as recruitment and funding issues. Consequently, HINT proves to be a reliable tool for identifying trial failures, particularly in later phases. Conversely, GPT-4o excelled in the early phases (such as Phase 1 and Phase 2), where it exhibited strong predictive capability and robustness, complementing the HINT model’s limitations in these stages.
@@
Large language models show distinct differences in predicting specific clinical trial endpoints. For example, they perform well on Overall Survival (OS) but struggle with endpoints like Objective Response Rate (ORR) and Progression-Free Survival (PFS). GPT-4o and GPT-4 demonstrate high accuracy in OS predictions; however, their performance on ORR and PFS often falls below random chance levels, indicating that these models are better suited for simpler survival endpoints. In contrast, the HINT model maintains more stable performance across various endpoints, with particular strength in recognizing negative samples.
The study also found that as trial duration increased, the predictive performance of the models generally declined. For extended trials, the HINT model’s predictive ability was especially limited. However, while GPT-4o and GPT-4 showed some performance decline, they managed to remain relatively stable.
@@
In conclusion, GPT-4o stands out as the best large language model for clinical trial prediction, but all LLMs face challenges in recognizing negative outcomes, particularly in complex areas like oncology. Nonetheless, these models are better suited for long-duration clinical trials. The HINT model’s high specificity in identifying negative outcomes highlights the need for further advancements in large language models to enhance their capability in negative predictions and in addressing external trial complexities.

### v2

@@ 2.4 Prediction for Different Clinical Endpoints Trials
The Fig.1 C illustrates the distribution of clinical trial outcomes across different model datasets(Table 4). Overall Survival (OS) dominates the outcome categories for all models, consistently representing around 64-70% of the data. Objective Response Rate (ORR) follows as the second most represented category, ranging from 6% to 11%. Progression-Free Survival (PFS) is included in a smaller proportion, around 3-4% across models. Safety indicators (AE and SAE) occupy a small portion, varying between 1.6% and 4.9%. Lastly, the ”Other” category (for example: Relapse-Free Survival, Pharmacokinetic Parameters, Patient-Reported Outcomes) accounts for around 14-19% of the outcomes across the datasets. Most clinical trials prioritize Overall Survival (OS) as the primary endpoint, and model performance on OS closely mirrors their balanced accuracy and MCC values.
@@
For other models, performance across different clinical trial endpoint categories showed similar trends. For instance, among all models that exhibited some recognition capability, the ability to identify clinical trials with Overall Survival (OS) as the endpoint was the strongest. For GPT-4o and GPT-4, balanced accuracy in this category was significantly above the average (0.538 and 0.523). This phenomenon was also observed in Llama3. Safety indicators (AE and SAE) also demonstrated recognition capabilities in these models. Although performance in this category was not as strong as in the OS category, metrics for GPT-4, GPT-4o, and Llama3 were better than random guess. In particular, GPT-4o’s specificity in this category reached 0.5, marking the highest level achieved by large language models in this experiment. However, the HINT model showed poor predictive performance for clinical trials focused on Overall Survival and safety indicators, with average accuracies of only 0.496 and 0.498, and MCC values of -0.028 and -0.007, respectively. This contrasts with the performance of the large models.
@@
For other clinical trial classifications, such as Objective Response Rate (ORR), Progression-Free Survival (PFS), and other uncategorized endpoints, most of these models did not exhibit classification capabilities. Metrics for these models remained at random guess levels. Notably, for GPT-4o, which typically performed well, the MCC for ORR was -0.167, even falling below 0. In contrast, the HINT model showed relatively better predictive performance on these two categories.
@@
Nonetheless, the specificity of the HINT model remained much higher than that of the large language models. This suggests that HINT maintained a relatively strong ability to recognize negative cases in these categories.
Endpoint | Trials Number | Balanced_acc | MCC | Recall | Specificity |
Model: GPT-4mini | |||||
Overall Survival (OS) | 28 | 0.500 | 0.0 | 1.000 | 0.0 |
Objective Response Rate (ORR) | 10 | 0.500 | 0.0 | 1.000 | 0.0 |
Progression-Free Survival (PFS) | 5 | 0.500 | 0.0 | 1.000 | 0.0 |
Safety Indicators (AE and SAE) | 12 | 0.500 | 0.0 | 1.000 | 0.0 |
Others | 13 | 0.000 | 0.0 | 1.000 | 0.0 |
Model: GPT-4o | |||||
Overall Survival (OS) | 28 | 0.614 | 0.278 | 1.000 | 0.333 |
Objective Response Rate (ORR) | 10 | 0.438 | -0.167 | 0.968 | 0.438 |
Progression-Free Survival (PFS) | 5 | 0.500 | 0.000 | 0.946 | 0.0 |
Safety Indicators (AE and SAE) | 12 | 0.583 | 0.169 | 0.962 | 0.500 |
Others | 13 | 0.517 | 0.000 | 0.962 | 0.158 |
Model: GPT-4 | |||||
Overall Survival (OS) | 20 | 0.562 | 0.281 | 1.000 | 0.125 |
Objective Response Rate (ORR) | 9 | 0.500 | 0.000 | 1.000 | 0.000 |
Progression-Free Survival (PFS) | 5 | 0.500 | 0.000 | 1.000 | 0.0 |
Safety Indicators (AE and SAE) | 9 | 0.600 | 0.316 | 1.000 | 0.200 |
Others | 7 | 1.000 | 0.000 | 1.000 | 0.062 |
Model: GPT-3.5 | |||||
Overall Survival (OS) | 1423 | 0.504 | 0.044 | 1.000 | 0.013 |
Objective Response Rate (ORR) | 409 | 0.499 | -0.023 | 0.998 | 0.000 |
Progression-Free Survival (PFS) | 163 | 0.506 | 0.040 | 0.998 | 0.029 |
Safety Indicators (AE and SAE) | 614 | 0.500 | -0.007 | 0.998 | 0.004 |
Others | 463 | 0.504 | 0.000 | 0.998 | 0.012 |
Model: HINT | |||||
Overall Survival (OS) | 252 | 0.548 | 0.085 | 0.596 | 0.515 |
Objective Response Rate (ORR) | 74 | 0.488 | -0.022 | 0.616 | 0.565 |
Progression-Free Survival (PFS) | 38 | 0.446 | -0.120 | 0.619 | 0.692 |
Safety Indicators (AE and SAE) | 89 | 0.516 | 0.034 | 0.605 | 0.366 |
Others | 67 | 0.511 | 0.000 | 0.605 | 0.566 |
Model: Llama3 | |||||
Overall Survival (OS) | 205 | 0.545 | 0.152 | 0.935 | 0.133 |
Objective Response Rate (ORR) | 61 | 0.507 | 0.018 | 0.952 | 0.111 |
Progression-Free Survival (PFS) | 26 | 0.476 | -0.098 | 0.951 | 0.000 |
Safety Indicators (AE and SAE) | 88 | 0.528 | 0.081 | 0.954 | 0.161 |
Others | 45 | 0.507 | 0.000 | 0.954 | 0.062 |
@@ 2.5 Prediction on different clinical trial period
We observe that for all models, the GPT-4mini and GPT-3.5 models showed poor (close to random) predictive ability in clinical trials of both short and long durations. In contrast, the GPT-4 and GPT-4o models generally demonstrated better predictive performance for medium- and long-term clinical trials. The GPT-4 model showed higher average accuracy (0.562) in trials lasting over 3000 days, but did not exhibit predictive power for shorter trials. GPT-4o showed optimal results for trials lasting between 1000 and 2000 days, but overall displayed an upward trend in performance. For the HINT and Llama3 models, results were similarly poor for trials lasting 1000-2000 days (average accuracies of 0.495 and 0.501, respectively), but both models showed good predictive ability for both short- and long-term trials.(Table 5).
When examining specificity, the HINT model shows an increasing trend over time, indicating an improved ability to recognize negative cases in longer trials. However, this comes at the cost of a decrease in recall. In contrast, the specificity trends of GPT-4, GPT-4o, and other large language models closely mirror the trends in their average accuracy, suggesting that the ability to recognize negative samples is closely tied to the overall predictive performance of these models.
This trend suggests that GPT-4 and GPT-4o models are more capable of predicting medium- and long-term clinical trials. In contrast, HINT and Llama3 perform better in predicting short-duration clinical trials. As the duration of the trial increases, the HINT model shows improved ability to recognize negative samples but at the expense of a decrease in its ability to identify positive samples. The other two more basic GPT models still do not demonstrate significant predictive capability for clinical trials.
Trial Period | Balanced Accuracy | MCC | Recall | Specificity |
Model: GPT-4mini | ||||
short-term trials | 0.500 | 0.000 | 1.000 | 0.000 |
medium-length trials | 0.500 | 0.000 | 1.000 | 0.000 |
long-term trials | 0.500 | 0.000 | 1.000 | 0.000 |
Model: GPT-4o | ||||
short-term trials | 0.417 | -0.218 | 0.833 | 0.000 |
medium-length trials | 0.575 | 0.200 | 0.900 | 0.250 |
long-term trials | 0.550 | 0.209 | 1.000 | 0.100 |
Model: GPT-4 | ||||
short-term trials | 0.500 | 0.000 | 1.000 | 0.000 |
medium-length trials | 0.500 | 0.000 | 1.000 | 0.000 |
long-term trials | 0.562 | 0.240 | 1.000 | 0.125 |
Model: GPT-3.5 | ||||
short-term trials | 0.504 | 0.058 | 0.998 | 0.011 |
medium-length trials | 0.503 | 0.043 | 0.998 | 0.008 |
long-term trials | 0.500 | 0.000 | 1.000 | 0.000 |
Model: HINT | ||||
short-term trials | 0.546 | 0.097 | 0.699 | 0.394 |
medium-length trials | 0.495 | -0.011 | 0.505 | 0.485 |
long-term trials | 0.524 | 0.049 | 0.364 | 0.684 |
Model: Llama3 | ||||
short-term trials | 0.560 | 0.248 | 0.985 | 0.137 |
medium-length trials | 0.501 | 0.007 | 0.951 | 0.052 |
long-term trials | 0.566 | 0.219 | 0.959 | 0.176 |
@@ 2.6 Terminated trials and prediction of termination
Large language models (such as the GPT series and Llama3) performed poorly in predicting outcomes for terminated trials, with accuracy around 0.1 or lower, indicating their difficulty in handling the complex external factors causing trial termination(Table 6). Consequently, we excluded “terminated” trials from further analyses to avoid biasing the experimental results. While large language models could not identify trial termination, the HINT model demonstrated this ability, with stable accuracy across all clinical trial statuses, similar to its performance on the entire dataset, showcasing its effectiveness in distinguishing terminated trials.
Model | HINT | GPT-3.5 | GPT-4o | GPT-4mini | GPT-4 | Llama3 |
@@ 3.3 Clinical Trials’ Label
The labeling process is carried out by two independent researchers using a double-blind cross-validation mechanism to ensure rigor. Discrepant cases are submitted for review by an independent panel consisting of clinical methodology experts and statisticians, who reach a final decision based on ICH[23] guidelines and statistical consensus. The core categories for clinical trial failure include the lack of statistical significance for the primary endpoint efficacy, failure to meet predefined efficacy targets or breach of non-inferiority boundaries, insufficient sample size or missing evaluable data leading to inadequate statistical power, absence of synergistic benefits in combination therapies, flaws in treatment cycle design that prevent efficacy from being observed at key time points, mid-term ineffectiveness analysis triggering early termination of the study, severe safety issues or changes in drug regulatory status leading to trial interruption, overall shift in efficacy trends without reaching the significance threshold, and failure to identify a clinically meaningful effective dose range during the dose-finding phase.
In subsequent analyses, we excluded data with a ” terminated ” status because we found that the LLMs lacked the ability to accurately recognize trial termination. Fig2 C illustrates the proportion of trial status labels across various trial statuses within the entire dataset.
@@ 3.4 Clinical Trials’ Prediction
For the HINT model, we input data in the required format[9], including the SMILES[24] sequence information of the drug and the ICD-10 code of the studied disease(Fig.2 E). Following the preprocessing steps outlined by Fu et al.[9], we categorized the data based on the trial phase before feeding it into the HINT model for prediction. The model output is a probability of trial success; results with a success probability greater than 0.5 were labeled as “Success,” while those with a probability less than or equal to 0.5 were labeled as “Failure”.
@@ 3.6 Metrics Selected
Recall (Sensitivity): The proportion of true positives correctly identified amonPT all actual positives. This metric measures the model’s ability to capture relevant instances and avoid false negatives.
@@
GPT-3.5 | gpt-3.5-turbo-0125 | Sep 2021 |
GPT-4 | gpt-4-turbo-2024-04-09 | Dec 2023 |
GPT-4o | chatgpt-4o-latest | Oct 2023 |
GPT-4mini | gpt-4o-mini-2024-07-18 | Oct 2023 |
Llama3 | Llama-3-8B-Instruct | March 2023 |
@@ 4 Conclusion
This study focuses on evaluating the performance of large language models (LLMs) in predicting clinical trial outcomes, with particular attention to Balanced Accuracy, Matthews Correlation Coefficient (MCC), Recall, and Specificity. Among all models, GPT-4o demonstrated the best overall performance, showing high balanced accuracy and MCC. However, similar to other LLMs, GPT-4o faced limitations in recognizing negative samples, a common challenge across these models. In comparison, the HINT model demonstrates consistently excellent ability to identify negative class samples.
@@
Different models show varying predictive abilities for clinical trials across different disease cateories. An important finding is that HINT struggles to predict the outcomes of oncology trials, which constitute the majority of the dataset. The complexity of cancer-related trials may be the reason for the model’s underperformance in this area compared to large language models. Despite HINT demonstrating advantages in predicting certain outcomes and maintaining accuracy in identifying negative samples, its limitations in handling complex fields like oncology restrict its overall predictive capacity. Llama3 performs the best in predicting oncology clinical trial outcomes, and GPT-4o also shows corresponding capabilities. Different large language models excel in predicting outcomes for different types of diseases.
In contrast, the HINT model shows relatively stable performance across different trial phases and exhibits a clear advantage in identifying negative samples. As the trial phase progresses, the effectiveness of the HINT model continues to improve. Furthermore, HINT is resilient to external factors such as recruitment and funding issues. In contrast, the predictive ability of large language models is less influenced by the trial phase. GPT-4o and Llama3 perform exceptionally well in early phases (such as Phase I and Phase II), demonstrating strong predictive capability and robustness, complementing the limitations of the HINT model in these phases. However, HINT maintains a stronger ability to identify negative samples than all large language models.
@@
Different models exhibit varying strengths in predicting specific clinical trial endpoints. Large language models, such as GPT-4o and GPT-4, show notable accuracy in predicting Overall Survival (OS), but struggle with more complex endpoints like Objective Response Rate (ORR) and Progression-Free Survival (PFS). Despite demonstrating strong performance in OS predictions, their ability to predict ORR and PFS often falls below random chance, indicating that they are more effective at handling simpler survival endpoints. In contrast, the HINT model demonstrates more consistent performance across various clinical trial endpoints and excels at recognizing negative samples, particularly in more complex categories such as ORR and PFS. While HINT’s predictive performance is not as strong for certain outcomes like OS, its superior specificity and ability to identify negative cases make it a valuable model in clinical trial prediction.
The study also found that as trial duration increased, the predictive performance of the models generally improved for medium- and long-term clinical trials. For extended trials, the HINT model demonstrated enhanced ability to recognize negative samples, although this came at the expense of decreased recall. Meanwhile, GPT-4o and GPT-4 showed relatively stable performance, with GPT-4 achieving higher accuracy in trials lasting over 3000 days
@@
In conclusion, GPT-4o and Llama3 stands out as the best large language model for clinical trial prediction, but all LLMs face challenges in recognizing negative outcomes.The performance of these two models varies in different scenarios and conditions, each with its own strengths and weaknesses. Nonetheless, these models are better suited for long-duration clinical trials. The HINT model’s high specificity in identifying negative outcomes highlights the need for further advancements in large language models to enhance their capability in negative predictions and in addressing external trial complexities.